6. Solve the following recurrence relations: (a) An+1 ,00 = 2 (b) n-1 an+1 =1+ ak , 0o = a1 = 1 ,n> 1 k=0
Answer #1

We know know C+D=1 C-l-D-1 - (SS-1) 25 C-255-55+ ✓st 255 255 -1 (1-5) 255 253 C= 1+55 an- c (145)+ 0 (1-5) nel ntl 1+5 2 5 (3